OpenStudy (anonymous):

Find the slope of the line (-2, 8) (1, -1)

OpenStudy (anonymous):

m=-3

OpenStudy (anonymous):

m=(y2-y1)/x2-x1) =(-1-8)/(1+2) = -9/3 m=-3

OpenStudy (anonymous):

thx